OpenStudy (anonymous):

Find the product of (x − 5)2. x2 + 10x + 25 x2 − 10x + 25 x2 − 25 x2 + 25

OpenStudy (hoblos):

(a-b)^2 = a2 - 2ab + b2 so (x − 5)2 = x2 − 10x + 25

jimthompson5910 (jim_thompson5910):

(x-5)^2 (x-5)(x-5) x*x+x*(-5)-5*x-5*(-5) x^2-5x-5x+25 x^2-10x+25 So (x-5)^2 expands out to x^2-10x+25 In other words, (x-5)^2=x^2-10x+25 for all values of x.